Operation Summary for June 18, 2010 - June 21, 2010 |
|
The Integrated Luminosity for the period from 6/14/10 to 6/21/10 was 58.92 pb-1. NuMI reported receiving 7.51E18 protons on target during this same period. Friday June 18 The day shift began with Operations monitoring Tevatron (TeV) store 7900 (142.76E30), and the Antiproton Source (Pbar) stack (76.2E10), with the Booster and the Main Injector (MI) and the Recycler in access, with the Linac operating normally, with the NuMI and MiniBooNE experiments waiting for beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, the MI access ended at 10 AM. The Booster access ended at 11 AM. Operations resumed sending beam to MiniBooNE, Pbar, and NuMI at 11:30 AM. Operations established beam to the Meson FTTB T-977 experiment at 12:30 PM. Preaccelerator personnel accessed the H- Source dome to check generator brushes at 1:15 PM. MI personnel conducted beam studies in the afternoon. A CUB Booster/Pbar 95 LCW pump failed at the end of shift; FESS personnel began repairs. During the evening shift, a lightning strike at 16:51 PM caused store 7900 to abort and a TeV main bus bypass (there was no quench). CDF and D0 accessed their halls for necessary repairs. The storm affected Booster and MI as well as tripping off various RF components. Operations prepared the TeV A-E and Transfer Hall for accesses to replace a failed vacuum turbo station in A-sector. Operations disabled the T-977 beam at 6:05 PM. MI personnel conducted beam studies. RF personnel investigated problems with a MI RF station (MIRF6). Operators conducted a TeV checkout and turn-on began at 9:30 PM; operators had the machine ready for a dry squeeze by the end of the shift. Operations contacted TeV personnel to investigate controls problems with the TeV sector B48 separator. Saturday June 19 The midnight shift began with Operations monitoring the stash (306.1E10), with Pbar stacking (8.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, TeV personnel continued working on the B48 separator; they swapped out the controller card and had the TeV ready for beam by 1:45 AM. Operations put the TeV into HEP shot setup and established store 7902 at 3:15 AM with an average initial luminosity (AIL) of 343.6E30. Operations put MIRF6 into standby due to an excessive number of repeated gap faults. During the day shift, Operations monitored systems. During the evening shift, Operations terminated store 7902 at 5:30 PM, put the TeV into shot setup, and established store 7903 at 7:30 PM with an AIL of 343.3E30. Sunday June 20 The midnight shift began with Operations monitoring store 7903 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported some minor downtime due to LRF5 and MIRF2. During the day shift, Operations terminated store 7903 at 10:10 AM. Operations put the TeV into shot setup, but a failed power supply at TeV sector B4 locked the main bus to 90 GeV. At noon, Operations reported that LRF5 tripped off several times and eventually failed to reset; operators called in Linac personnel and they replaced a bad keyer module in the modulator. During the evening shift, Linac resumed sending beam at 4:45 PM. A TeV E3 sump pump failed; operators prepared F-Sector for access to install a portable sump pump. Operations had the TeV recovery underway by 5 PM. Operators conducted a dry squeeze and checked the orbits. Operations put the TeV into shot setup at 7:45 PM and established store 7908 at 9:05 PM with an AIL of 331E30. Monday June 21 The midnight shift began with Operations monitoring store 7908 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-977 at 4:40 AM. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 06/18/2010 email Fermilab |
| Security, Privacy, Legal | |